News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 11 August 2000 1.00pm EDT

During the last week, Chandra completed the observing schedule as planned.

A real-time procedure to insert the LETG was conducted at 12.30am EDT on 7 Aug and was completed as planned with nominal telemetry indicators. The grating was inserted to perform a block of planned LETG observations during last weeks daily load. A real-time procedure is planned for 6.20pm EDT on 12 Aug to retract the LETG.

Other real-time activities planned for the coming week include running the 9th ACIS Squeegee mode test at 6.30am EDT on Aug 12, and an on-orbit demonstration of a procedure to off-point the solar arrays during a lunar eclipse that will occur in October. The off-pointing is required to avoid a power bus over-voltage condition at the completion of the eclipse. The demonstration scheduled for Aug 12 will determine the correct solar array angle for use during the eclipse in October.

Also of note last week was the successful completion of the AO2 peer review. 186 winning proposals were selected from 732 submitted and announcement letters are expected to be sent next week as planned.

The scheduled observations for the next week are:

------------------------------------------
G320.4-1.2              ACIS-I       Aug 14
PSR B1706-44            ACIS-S
SGR 1806-20             ACIS-S       Aug 15
2EG J1811-2339          ACIS-I
NGC 253                 ACIS-S       Aug 16
NGC 1407                ACIS-S
MITG J0414+0534         ACIS-S
MBM 12                  ACIS-S
Radiation Belt                       Aug 17
CTI                     ACIS         Aug 18
M31 (multiple)          HRC-I
MS 0015.9+1609          ACIS-I
A85                     ACIS-I       Aug 19
AWM 7                   ACIS-I
Radiation Belt                       Aug 20
0235+164                ACIS-S

------------------------------------------

The spacecraft subsystems continued to operate nominally.

Return to Status Reports